if the distance between two positive point charges is tripled,then the strength of the electrostatic repulsion between them will decrease by a factor of??

Grade:

3 Answers

Chetan Mandayam Nayakar
12 years ago

the strength of repulsion is proportional to 1/r2, therefore is r is tripled repulsion strength will decrease by a factor of 9

Gaurav Yadav
12 years ago

Force 1 (F1) = kqXq/rXr

Force 2 (F2) = kqXq/9rXr      ( now r will become 3r )

F2/F1 = 1/9

so we can say that if distance between to positive charges is tripled then the electrostatic repulsion between them will become 1/9 or one ninth of the orignal value.

 

Please approve my answer if you liked it by clicking on "Yes" given below..!!! 

anapurna monga
12 years ago

Cool9